if the ratio of magnetic field at two points in a difinte direction due to a long current carrying wire is ¾  then the ratio of the distances of these points from the wire will be??
give solution

Grade:12

1 Answers

Demo_Gulbarga
4 years ago
B = mu_0  I / (2 pi r)  ;  by ampere’s circuital law
                                  ; amperean loop is circle of radius r, with centre at wire location and area of circle perpendicular to wire
 
B1 / B2 = r2 / r1 = 3 / 4
==> r1 / r2 = 4 / 3